# Ulrich Bundles on Jacobian Variety of a Curve

### 04 Mar 2026 v2

@@ Abstract
Let $C$ be a smooth complex projective curve of genus $g>1$ and $A=J(C)$ its Jacobian with principal polarization $\Theta$. Starting from a semistable vector bundle $V$ on $C$ with $\mu(V)>2g-2$, we consider the Fourier–Mukai transform $E=\Phi_{\mathcal{P}}(a_{*}V)$. We prove that $E(\Theta)$ satisfies the $\mathrm{IT}_{0}$ property. As a consequence, for every $m\geq g+1$, the polarized Jacobian $(A,m\Theta)$ admits Ulrich bundles constructed functorially from $V$. Further we analyze stability and Chern classes of the resulting bundles. We show that the construction induces a natural morphism from the generically finite cover of moduli space of stable bundles on $C$ to the moduli space of stable bundles on $A$, producing positive-dimensional families of stable Ulrich bundles.
@@ 1 Introduction
Ulrich bundles occupy a central position in the interaction between algebraic geometry and commutative algebra. Originally introduced by Eisenbud and Schreyer in connection with linear determinantal representations and maximal Cohen–Macaulay modules [3]. Ulrich bundles are vector bundles charaterized by strong cohomological vanishing properties with respect to given polarization. they play a central role in study of syzygies and linear resolutions of projective varieties and Chow forms of projective varieties.
A fundamental conjecture of Eisenbud and Schreyer predicts that every polarized projective variety admits an Ulrich bundle. Despite substantial progress in special cases, the conjecture remains widely open. Ulrich bundles have been constructed for several classes of varieties, including many hypersurfaces and certain surfaces; see for example [1, 7]. However, the existence problem remains largely unexplored for irregular varieties and higher-dimensional abelian varieties.
@@
In this paper we study the existence of Ulrich bundles on abelian varieties arising from Jacobians of smooth projective curves. Let $C$ be a smooth projective curve of genus $g\geq 2$, and let $A=J(C)$ denote its Jacobian equipped with the principal polarization $\Theta$. Our approach is based on Fourier–Mukai techniques and exploits the geometry of the Abel–Jacobi embedding $C\hookrightarrow J(C)$. Starting from a sufficiently positive line bundle $L$ on $C$, we consider the pushforward $a_{*}L$ along the Abel–Jacobi map and study its Fourier–Mukai transform. After a suitable twist by the principal polarization, we show that the resulting bundle on $J(C)$ satisfies strong regularity properties in the sense of Pareschi–Popa [8]. Using the theory of continuously globally generated bundles on abelian varieties, we construct a kernel bundle whose twist produces an Ulrich bundle on $(J(C),m\Theta)$ for sufficiently large $m$.
Our main result as follows.
Theorem A (See Theorem 5.2). Let $C$ be a smooth projective curve of genus $g>1$ and let $V$ be a semistable vector bundle on $C$ with $\mu(V)>2g-2$. Then for every integer $m\geq g+1$, the polarized Jacobian $(J(C),m\Theta)$ admits an Ulrich bundle constructed functorially from $V$.
This construction verifies the Eisenbud–Schreyer conjecture for a large and geometrically natural class of abelian varieties, namely Jacobians of smooth projective curves endowed with sufficiently positive multiples of the principal polarization. To the best of our knowledge, this provides the first systematic construction of Ulrich bundles on Jacobians in arbitrary genus with respect to sufficiently positive polarization. It may be viewed as a higher-dimensional analogue of Beauville’s construction of Ulrich bundles on abelian surfaces [2].
Beyond the existence statement, the construction has further geometric consequences. The Ulrich bundles obtained here arise naturally from vector bundles associated to the curve and vary in positive-dimensional families. More precisely, the construction induces families of stable bundles on the Jacobian arising from moduli spaces of vector bundles on the curve.
Theorem B (See Theorem 10.2)
Fix integers $r\geq 2,d>r(2g-2)$, and let $M_{C}$ denote the moduli space of stable vector bundles on $C$ of rank $r$ and degree $d$. For $m\geq g+1$, let $M_{A}$ be the moduli space of $H$-stable bundles on $(A,m\Theta)$ with the numerical invariants of the Ulrich bundles constructed in Theorem 5.2 Then there exists a quasi-projective variety $T$, a generically finite surjective morphism
$\pi:T\longrightarrow M_{C},$
and a dense open subset $T^{\mathrm{st}}\subset T$ such that the Ulrich construction is compatible with families and induces a morphism
$\Phi:T^{\mathrm{st}}\longrightarrow M_{A}.$
In particular, $M_{A}$ contains positive-dimensional families of stable Ulrich bundles parametrized by a variety of dimension $(r^{2}-1)(g-1)$.
These results illustrate a close relationship between the geometry of the curve and the moduli theory of vector bundles on its Jacobian. In particular, the construction connects Fourier–Mukai transforms, the geometry of Picard bundles, and the theory of Ulrich bundles. The techniques used here rely heavily on the modern theory of regularity on abelian varieties developed by Pareschi and Popa [8]. Their results on $IT_{0}$ sheaves and continuously globally generated bundles provide the key tools for controlling the cohomological behaviour of the Fourier–Mukai transforms appearing in our construction. We also make use of Mukai’s theory of semihomogeneous vector bundles [16].
Layout of the paper. In Section 2 we review some of the pre-requisites : Ulrich bundles, Abel-Jacobi map,Fourier–Mukai transforms, IT index properties, stable bundles and, semi-homogeneous bundles. We establish the $\mathrm{WIT}_{0}$ property and computes the rank of the transform in Section 3. In Section 4 we prove that $E(\Theta)$ satisfies $\mathrm{IT}_{0}$. Section 5 constructs Ulrich bundles and proves Theorem A. In Section 6 we discuss the numerical bound $m\geq g+1$. Section 7 studies slope stability. We compute Chern classes in Section 8. In Section 9 we specialize to line bundles and derive an upper bound for Ulrich complexity. We conclude Section 10 with the construction of the moduli morphism and proves the Theorem B.
Throughout the paper we work over the field $\mathbb{C}$.
2 Preliminaries
2.1 Ulrich bundles and Ulrich complexity
Let $(X,H)$ be a polarized smooth projective variety of dimension $n$, where $H$ is an ample divisor on $X$. A vector bundle $E$ on $X$ is called an Ulrich bundle with respect to $H$ if
$H^{i}(X,E(-jH))=0\qquad\text{for all }i\text{ and }1\leq j\leq n.$
In this paper we will consider Ulrich bundles on polarized Jacobians
$(A,m\Theta),$
where $A=J(C)$ is the Jacobian of a smooth complex projective curve $C$ of genus $g>1$ and $\Theta$ denotes the principal polarization.
Ulrich complexity. For a polarized variety $(X,H)$ admitting Ulrich bundles, we define the Ulrich complexity to be
$\mathrm{uc}(X,H):=\min\{rk(E)\;:\;E\text{ is an Ulrich bundle on }(X,H)\}.$
In general, determining $\mathrm{uc}(X,H)$ is a subtle problem. In the present work, we construct explicit Ulrich bundles on $(J(C),m\Theta)$ and derive upper bounds for $\mathrm{uc}(J(C),m\Theta)$ in terms of the genus $g$ and the degree of the initial bundle on $C$.
@@ 2.2 The Abel–Jacobi embedding
is a closed immersion. We identify $C$ with its image in $A$.
A classical result (see [5, Ch. VII]) states that
@@ 2.3 Fourier–Mukai transform on the Jacobian
Let $\widehat{A}:=Pic^{0}(A)$ be the dual abelian variety. Since $A$ is principally polarized, we identify
@@ 2.4 WIT and IT conditions
Let $\mathcal{F}$ be a coherent sheaf on $A$.
Definition 2.1.
We say that $\mathcal{F}$ satisfies WITi (Weak Index Theorem of index $i$) if
@@
A coherent sheaf $\mathcal{E}$ on $A$ satisfies IT0 (Index Theorem with index 0) if
@@ 2.5 Base change for Fourier–Mukai
Let $\mathcal{F}$ be a coherent sheaf on $A$. For $\alpha\in Pic^{0}(A)$, derived base change gives
@@ 2.6 Slope Stability
2.7 Semihomogeneous bundles on abelian varieties
Let $A$ be an abelian variety and denote by $t_{x}:A\to A$ the translation by a point $x\in A$.
Definition 2.4.
A vector bundle $\mathcal{E}$ on $A$ is called semihomogeneous if for every point $x\in A$ there exists a line bundle $\alpha_{x}\in Pic^{0}(A)$ such that
$t_{x}^{*}\mathcal{E}\cong\mathcal{E}\otimes\alpha_{x}.$
Semihomogeneous bundles were introduced and studied by Mukai [16]. They enjoy the following properties:
  • A semihomogeneous bundle is simple if and only if it is stable with respect to any polarization [16, Prop. 6.16].
  • If $\mathcal{E}$ is semihomogeneous, then its Chern classes are proportional to powers of $c_{1}(\mathcal{E})$ in the Néron–Severi group [16, Thm. 5.8].
@@ 3.2 Derived fibers and WIT0
    $E$ is a locally free sheaf on $A$;
@@ 3.3 Rank formula
The result follows from Proposition $3.$ ∎
@@ 4.4 Vanishing of higher cohomology
Since $C$ has dimension $1$, we have

### 06 Mar 2026 v3

# Pabitra Barik

@@ Abstract
Let $C$ be a smooth complex projective curve of genus $g>1$ and $A=J(C)$ its Jacobian with principal polarization $\Theta$. Starting from a semistable vector bundle $V$ on $C$ with $\mu(V)>2g-2$, we consider the Fourier–Mukai transform $E=\Phi_{\mathcal{P}}(a_{*}V)$. We prove that $E(\Theta)$ satisfies the $\mathrm{IT}_{0}$ property.
@@ 1 Introduction
Let $C$ be a smooth projective curve of genus $g\geq 2$, and let $A=J(C)$ denote its Jacobian equipped with the principal polarization $\Theta$. Our approach is based on Fourier–Mukai techniques and exploits the geometry of the Abel–Jacobi embedding $C\hookrightarrow J(C)$. Starting from a sufficiently positive semistable vector bundle bundle $V$ on $C$, we consider the pushforward $a_{*}V$ along the Abel–Jacobi map and study its Fourier–Mukai transform. After a suitable twist by the principal polarization, we show that the resulting bundle on $J(C)$ satisfies strong regularity properties in the sense of Pareschi–Popa [4].
We recall some preliminaries.
@@ 1.1 The Abel–Jacobi embedding
is a closed immersion. We identify $C$ with its image in AA.
A classical result (see [2, Ch. VII]) states that
@@ 1.2 Fourier–Mukai transform on the Jacobian
Let $\widehat{A}:=Pic^{0}(A)$ be the dual abelian variety. Since AA is principally polarized, we identify
@@ 1.3 WIT and IT conditions
Let $\mathcal{F}$ be a coherent sheaf on AA.
Definition $1.$
We say that $\mathcal{F}$ satisfies WITi (Weak Index Theorem of index ii) if
@@
A coherent sheaf $\mathcal{E}$ on AA satisfies IT0 (Index Theorem with index 0) if
@@ 1.4 Base change for Fourier–Mukai
Let $\mathcal{F}$ be a coherent sheaf on AA. For $\alpha\in Pic^{0}(A)$, derived base change gives
@@ 2.2 Derived fibers and WIT0
    $E$ is a locally free sheaf on AA;
@@ 2.3 Rank formula
The result follows from Proposition 2.3. ∎
@@ 4.4 Vanishing of higher cohomology
Since $C$ has dimension 11, we have